# Broadening of the Divertor Heat Flux Profile in High Confinement Tokamak Fusion Plasmas with Edge Pedestals Limited by Turbulence in DIII-D

### v5

@@ Abstract
Figure 1: (a) Measured $\lambda_{q}$ from Langmuir probes (LP) as a function of BES measured high-frequency electron mode density fluctuation amplitude $\delta n/n$ at $\rho=0.97$, and (b) Corresponding measured divertor leg diffusive spreading $S$. Results from XGC simulations are shown in black open symbols using measured values from No. 184833 for the $x$-coordinate (noise prevents determination of $S$ from XGC for WPQH). Error bars represent statistical standard deviations in the time averages. Definitions of $\lambda_{q}$ and $S$ are given in Appendix A.
@@
In this Letter we demonstrate experimentally that the divertor heat flux width (referred to the outer midplane) $\lambda_{q}$ increases with the intensity of high frequency turbulence propagating in the electron diamagnetic direction near the last closed flux surface. We show that for quiescent H-mode plasmas in the DIII-D tokamak where edge turbulence is sufficiently strong, $\lambda_{q}$ does not follow the empirical scaling, and can even increase favorably with poloidal magnetic field. In these experiments we have used actuators (varying applied toroidal torque and plasma current) to control the edge turbulence. Multiple diagnostic measurements of both edge turbulence and divertor heat flux profiles show that as the electron turbulence intensifies, the fraction of the divertor heat flux carried by electrons increases and its profile broadens, broadening the total heat flux profile. The endpoints of the measured $\lambda_{q}$ range are quantitatively matched within measurement uncertainty by XGC gyrokinetic particle simulations, as summarized in Figs. 1 and 2. The turbulence produces large density fluctuations measured by beam emission spectroscopy (BES) [McKee:1999]. Both $\lambda_{q}$ and $S$ increase with measured density fluctuation amplitude, so that the total integral heat flux width $\lambda_{\mathrm{int}}\approx\lambda_{q}+1.64\,S$ also increases, as discussed in Appendix A. The mechanism for the broadening identified in these experiments lends plausibility to analogous XGC predictions of $\lambda_{q}\sim 6$ mm for ITER [chang:2017, chang:2021].
@@
Figure 3 shows the temporal evolution for the DIII-D discharge No. 184833 exhibiting the greatest $\lambda_{q}$ broadening, shown in Fig. 1. We first establish an H-mode edge pedestal limited by turbulent transport without ELMs, exhibiting a high and wide pedestal pressure profile, referred to as the wide pedestal quiescent H-mode (WPQH-mode) regime [burrell:2016, ernst:2018, burrell:2020, chen:2020]. Its pedestal width in normalized poloidal magnetic flux significantly exceeds that predicted by EPED scaling [snyder:2011, snyder:2012], given by $0.089\beta_{p}^{1/2}$, where $\beta_{p}=8\pi p/B_{p}^{2}$ is the pedestal top ratio of kinetic pressure to magnetic pressure, as shown. This regime is accessed through wall conditioning (boronization) to reduce collisionality, with low neutral beam injected (NBI) toroidal torque <2<2 Nm. During the period 4-5 s the counter-$I_{p}$ NBI torque is increased from 0.6 to 4 Nm, provoking a back transition to a little explored variant of standard quiescent H-mode (QH-mode) which also exhibits only broadband pedestal turbulence [burrell:2005], referred to here as “turbulent QH-mode.” More typical standard QH-mode pedestals are limited by low toroidal mode number N=15N=1-5 Edge Harmonic Oscillations (EHOs). The energy confinement time, pedestal pressure, and pedestal width step down significantly at this transition, indicating increased turbulent transport. Pedestal radial profiles are shown in Appendix B. The lower divertor heat flux profiles are measured by fixed Langmuir probes and infrared (IR) camera imaging using strike point sweeps in each phase. Close agreement between Langmuir probes and IR thermography is found when using the standard sheath heat flux transmission coefficient to relate total parallel heat flux to Langmuir probe measured electron saturation current and electron temperature [Stangeby:2000]. The comparison, together with detailed measurements for all discharges is given in Appendix C.
@@
Both the WPQH-mode and QH-mode phases of discharge No. 184833 are simulated with the XGC total-$f$ gyrokinetic particle code [chang:2017], using gyrokinetic ions and drift-kinetic electrons, and including Monte Carlo neutrals with recycling coefficient 0.99, the carbon impurity density as measured, and electromagnetic effects. Radial pedestal profiles used as input to XGC in are shown in Appendix B. Focusing on the electron mode turbulence, toroidal mode numbers N=15N=1-5 are filtered out to avoid possible MHD related fast instabilities (this may artificially reduce the lowest frequency ion-directed turbulence amplitude). The overall results for $\lambda_{q}$ are shown in Fig. 2, overlayed on the multimachine database [eich:2013].
@@
Finally, scaling arguments suggest that in future machines, pedestal turbulent transport may be sufficiently limiting to prevent ELMs [ernst:2018, ernst:2022aps], so that non-ELMing regimes naturally arise (see Appendix D for details). Intrinsically non-ELMing operating regimes, such as wide pedestal quiescent H-mode and quiescent H-mode, offer a high-performance solution at relevant low pedestal top collisionalities, featuring divertor heat flux profiles broadened by turbulence.
@@ Acknowledgements.
Appendix A: Eich Function.— The integral heat flux profile width $\lambda_{\mathrm{int}}$ can be related to the Eich fit parameters $(\lambda_{q},S)$ by $\lambda_{\mathrm{int}}=\int(q_{\|}(s)-q_{\mathrm{BG}})ds/(q_{\|0})\approx\lambda_{q}+1.64\,S$ [makowski:2012]. The Eich function is given by
@@
Appendix B: Profile Fits.— Pedestal profiles, shown in Fig. 9, are constrained by Langmuir probe measurements of separatrix electron temperature $T_{e}^{\mathrm{s}ep}$ and pedestal carbon and deuterium charge exchange measurements, which show constant carbon concentration across the pedestal. Equilibrium reconstructions of discharge No. 184833 in WPQH and QH phases, including the bootstrap current and realistic heating profiles from interpretative transport simulations (TRANSP) are used as input to the XGC code.
Appendix C: Divertor Heat Flux Measurements.— Using the standard sheath heat transmission coefficient $\gamma_{\mathrm{SH}}=7$ [Stangeby:2000] to relate total parallel heat flux from electrons and ions to Langmuir probe measured saturation current ($J_{\mathrm{sat}}$) and electron temperature ($T_{e}$) via $q_{\|}^{\mathrm{tot}}=\gamma_{\mathrm{SH}}J_{\mathrm{sat}}T_{e}/e$, close agreement is obtained between Langmuir probe measured divertor heat flux profiles and IR thermography measured heat flux profiles, as shown in Fig. 11. The IR measurements are shown for corresponding cases where available. This agreement translates to agreement within statistical measurement uncertainty in the $\lambda_{q}$ values inferred from Langmuir probes and IR, as shown in Fig. 10. Importantly, the $\lambda_{q}$ inferred from Langmuir probes is independent of the assumed $\gamma_{\mathrm{SH}}$. Because Langmuir probes measure mainly electrons, this result can be viewed as evidence that the electron contribution to the divertor heat flux underlies the broadening of $\lambda_{q}$ measured by IR.
@@
Appendix D: Scaling Pedestal Turbulent Transport to Future Machines.— Here we provide details on the scaling of pedestal turbulent transport with ion gyroradius, which varies inversely with magnetic field. Turbulent transport reduction by sheared flows is predicted to scale with the normalized ion gyroradius, $\rho_{*}=\rho_{i}/a$, where $a$ is the plasma minor radius [kotschenreuther:1996]. In future machines, $\rho_{*}$ is expected to be approximately 3 times smaller than in present tokamaks. The radial electric field $E_{r}$ in the edge pedestal is $Zen_{i}E_{r}\sim dp_{i}/dr\sim p_{i}/\Delta$, where $Zen_{i}$ is the ion charge density, $p_{i}=n_{i}T_{i}$ is the ion pressure with $T_{i}$ the ion temperature, $r$ is the minor radius, and $\Delta$ is the gradient scale length The $E\times B$} shear rate for turbulence is then $\gamma_{E}\sim B^{-1}dE_{r}/dr\sim E_{r}/(B\Delta)$ where $B$ is the magnetic field. Estimating the growth rate for ion scale drift-type instabilities as $\gamma_{\mathrm{lin}}\sim v_{\mathrm{thi}}/\Delta$, where $v_{\mathrm{thi}}=(2T_{i}/m_{i})^{1/2}$ is the ion thermal speed, the reduction of turbulent transport by $E\times B$} shear will depend on a parameter which scales as $\gamma_{E}/\gamma_{\mathrm{lin}}\sim\rho_{i}/\Delta\sim(a/\Delta)\rho_{*}$. The ratio of pedestal turbulent heat flux $Q$ to the gyroBohm flux $Q_{\mathrm{gB}}$ is predicted to increase asymptotically as a result of the decreased $E\times B$} shear at low $\rho_{*}$ according to [zhang:1992, kotschenreuther:2017, hatch:2018]

### v6

@@ Abstract
Figure 1: (a) Measured $\lambda_{q}$ from Langmuir probes (LP) as a function of BES measured high-frequency electron mode density fluctuation amplitude $\delta n/n$ at $\rho=0.97$, and (b) Corresponding measured divertor leg diffusive spreading $S$. Results from XGC simulations are shown in black open symbols using measured values from No. 184833 for the $x$-coordinate (noise prevents determination of $S$ from XGC for WPQH). Error bars represent statistical standard deviations in the time averages. Definitions of $\lambda_{q}$ and $S$ are given in the Appendix.
@@
In this Letter we demonstrate experimentally that the divertor heat flux width (referred to the outer midplane) $\lambda_{q}$ increases with the intensity of high frequency turbulence propagating in the electron diamagnetic direction near the last closed flux surface. We show that for quiescent H-mode plasmas in the DIII-D tokamak where edge turbulence is sufficiently strong, $\lambda_{q}$ does not follow the empirical scaling, and can even increase favorably with poloidal magnetic field. In these experiments we have used actuators (varying applied toroidal torque and plasma current) to control the edge turbulence. Multiple diagnostic measurements of both edge turbulence and divertor heat flux profiles show that as the electron turbulence intensifies, the fraction of the divertor heat flux carried by electrons increases and its profile broadens, broadening the total heat flux profile. The endpoints of the measured $\lambda_{q}$ range are quantitatively matched within measurement uncertainty by XGC gyrokinetic particle simulations, as summarized in Figs. 1 and 2. The turbulence produces large density fluctuations measured by beam emission spectroscopy (BES) [McKee:1999]. Both $\lambda_{q}$ and $S$ increase with measured density fluctuation amplitude, so that the total integral heat flux width $\lambda_{\mathrm{int}}\approx\lambda_{q}+1.64\,S$ also increases, as discussed in the Appendix. The mechanism for the broadening identified in these experiments lends plausibility to analogous XGC predictions of $\lambda_{q}\sim 6$ mm for ITER [chang:2017, chang:2021].
@@
Figure 3 shows the temporal evolution for the DIII-D discharge No. 184833 exhibiting the greatest $\lambda_{q}$ broadening, shown in Fig. 1. We first establish an H-mode edge pedestal limited by turbulent transport without ELMs, exhibiting a high and wide pedestal pressure profile, referred to as the wide pedestal quiescent H-mode (WPQH-mode) regime [burrell:2016, ernst:2018, burrell:2020, chen:2020]. Its pedestal width in normalized poloidal magnetic flux significantly exceeds that predicted by EPED scaling [snyder:2011, snyder:2012], given by $0.089\beta_{p}^{1/2}$, where $\beta_{p}=8\pi p/B_{p}^{2}$ is the pedestal top ratio of kinetic pressure to magnetic pressure, as shown. This regime is accessed through wall conditioning (boronization) to reduce collisionality, with low neutral beam injected (NBI) toroidal torque <2<2 Nm. During the period 4-5 s the counter-$I_{p}$ NBI torque is increased from 0.6 to 4 Nm, provoking a back transition to a little explored variant of standard quiescent H-mode (QH-mode) which also exhibits only broadband pedestal turbulence [burrell:2005], referred to here as “turbulent QH-mode.” More typical standard QH-mode pedestals are limited by low toroidal mode number N=15N=1-5 Edge Harmonic Oscillations (EHOs). The energy confinement time, pedestal pressure, and pedestal width step down significantly at this transition, indicating increased turbulent transport. Pedestal radial profiles are shown in the Appendix. The lower divertor heat flux profiles are measured by fixed Langmuir probes and infrared (IR) camera imaging using strike point sweeps in each phase. Close agreement between Langmuir probes and IR thermography is found when using the standard sheath heat flux transmission coefficient to relate total parallel heat flux to Langmuir probe measured electron saturation current and electron temperature [Stangeby:2000]. The comparison, together with detailed measurements for all discharges is given in the Appendix.
@@
Both the WPQH-mode and QH-mode phases of discharge No. 184833 are simulated with the XGC total-$f$ gyrokinetic particle code [chang:2017], using gyrokinetic ions and drift-kinetic electrons, and including Monte Carlo neutrals with recycling coefficient 0.99, the carbon impurity density as measured, and electromagnetic effects. Radial pedestal profiles used as input to XGC in are shown in the Appendix. Focusing on the electron mode turbulence, toroidal mode numbers N=15N=1-5 are filtered out to avoid possible MHD related fast instabilities (this may artificially reduce the lowest frequency ion-directed turbulence amplitude). The overall results for $\lambda_{q}$ are shown in Fig. 2, overlayed on the multimachine database [eich:2013].
@@
Finally, scaling arguments suggest that in future machines, pedestal turbulent transport may be sufficiently limiting to prevent ELMs [ernst:2018, ernst:2022aps], so that non-ELMing regimes naturally arise (see the Appendix for details). Intrinsically non-ELMing operating regimes, such as wide pedestal quiescent H-mode and quiescent H-mode, offer a high-performance solution at relevant low pedestal top collisionalities, featuring divertor heat flux profiles broadened by turbulence.
@@ Acknowledgements.
* Eich Function.— The integral heat flux profile width $\lambda_{\mathrm{int}}$ can be related to the Eich fit parameters $(\lambda_{q},S)$ by $\lambda_{\mathrm{int}}=\int(q_{\|}(s)-q_{\mathrm{BG}})ds/(q_{\|0})\approx\lambda_{q}+1.64\,S$ [makowski:2012]. The Eich function is given by
@@
Profile Fits.— Pedestal profiles, shown in Fig. 9, are constrained by Langmuir probe measurements of separatrix electron temperature $T_{e}^{\mathrm{s}ep}$ and pedestal carbon and deuterium charge exchange measurements, which show constant carbon concentration across the pedestal. Equilibrium reconstructions of discharge No. 184833 in WPQH and QH phases, including the bootstrap current and realistic heating profiles from interpretative transport simulations (TRANSP) are used as input to the XGC code.
Divertor Heat Flux Measurements.— Using the standard sheath heat transmission coefficient $\gamma_{\mathrm{SH}}=7$ [Stangeby:2000] to relate total parallel heat flux from electrons and ions to Langmuir probe measured saturation current ($J_{\mathrm{sat}}$) and electron temperature ($T_{e}$) via $q_{\|}^{\mathrm{tot}}=\gamma_{\mathrm{SH}}J_{\mathrm{sat}}T_{e}/e$, close agreement is obtained between Langmuir probe measured divertor heat flux profiles and IR thermography measured heat flux profiles, as shown in Fig. 11. The IR measurements are shown for corresponding cases where available. This agreement translates to agreement within statistical measurement uncertainty in the $\lambda_{q}$ values inferred from Langmuir probes and IR, as shown in Fig. 10. Importantly, the $\lambda_{q}$ inferred from Langmuir probes is independent of the assumed $\gamma_{\mathrm{SH}}$. Because Langmuir probes measure mainly electrons, this result can be viewed as evidence that the electron contribution to the divertor heat flux underlies the broadening of $\lambda_{q}$ measured by IR.
@@
Scaling Pedestal Turbulent Transport to Future Machines.— Here we provide details on the scaling of pedestal turbulent transport with ion gyroradius, which varies inversely with magnetic field. Turbulent transport reduction by sheared flows is predicted to scale with the normalized ion gyroradius, $\rho_{*}=\rho_{i}/a$, where $a$ is the plasma minor radius [kotschenreuther:1996]. In future machines, $\rho_{*}$ is expected to be approximately 3 times smaller than in present tokamaks. The radial electric field $E_{r}$ in the edge pedestal is $Zen_{i}E_{r}\sim dp_{i}/dr\sim p_{i}/\Delta$, where $Zen_{i}$ is the ion charge density, $p_{i}=n_{i}T_{i}$ is the ion pressure with $T_{i}$ the ion temperature, $r$ is the minor radius, and $\Delta$ is the gradient scale length The $E\times B$} shear rate for turbulence is then $\gamma_{E}\sim B^{-1}dE_{r}/dr\sim E_{r}/(B\Delta)$ where $B$ is the magnetic field. Estimating the growth rate for ion scale drift-type instabilities as $\gamma_{\mathrm{lin}}\sim v_{\mathrm{thi}}/\Delta$, where $v_{\mathrm{thi}}=(2T_{i}/m_{i})^{1/2}$ is the ion thermal speed, the reduction of turbulent transport by $E\times B$} shear will depend on a parameter which scales as $\gamma_{E}/\gamma_{\mathrm{lin}}\sim\rho_{i}/\Delta\sim(a/\Delta)\rho_{*}$. The ratio of pedestal turbulent heat flux $Q$ to the gyroBohm flux $Q_{\mathrm{gB}}$ is predicted to increase asymptotically as a result of the decreased $E\times B$} shear at low $\rho_{*}$ according to [zhang:1992, kotschenreuther:2017, hatch:2018]